Could you give me the details of solving this:Given the following information and assuming a CCA rate of 20%, what is the NPV of this project? Initial investment = $400,000; life = five years; before-tax cost savings = $150,000 per year; salvage value = $30,000 in year 5; tax rate = 34%; discount rate = 14%. A. -$149,841 B. -$33,117 C. $0 D. $19,800 E. $27,428
Answer #1

NPV = pv of inflow - pv of outflow

Pv of inflow = (150000-34%)*pvaf(14%,5) + (30000-34%)*pvf(14%5)

= 99000*3.433 + 19800*.5198

=339867 + 10292

= 350159

Pv of outflow = initial investment = 400000

Hence, npv = 400000-350159 = 49841